Find the stock of money in an economy where the velocity of money is 3, long run real output is 700,000 units and the long run price is $6.
1
Expert's answer
2019-06-06T08:56:25-0400

The formula is: MV = PQ.

If the velocity of money is 3, long run real output is 700,000 units and the long run price is $6, then:

M = PQ/V = 6×700,000/3 = $1,400,000.